HUBBLE SPACE TELESCOPE DAILY REPORT #1840 PERIOD COVERED: 0000Z (UTC) 03/21/97 - 0000Z (UTC) 03/24/97 Daily Status Report as of 083/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7092 (CCD Dark Rate and Read Noise) Observations were taken to measure CCD dark current and gather = data for making super darks. The observations were completed as planned, = and no anomalies were reported. 1.2 Completed STIS/CCD 7063 (STIS SMOV CCD Contamination Monitor) The Space Telescope Imaging Spectrograph CCD was used to make = observations of standard star G191B2B in order to measure flux and look = for any gross changes over time that might indicate contamination. The = observations were completed as scheduled, and no anomalies were reported. = 1.3 Completed Two Sets of WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: = Decontamination) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 1.4 Completed Two Sets of WF/PC-2 6361 (High Resolution Imaging of = X-Ray Selected AGNs) The WF/PC-2 was used to make high resolution observations of the = known, low, red-shifted, x-ray active galactic nuclei (AGN) MS0818.8+5 = and MS1112.5+4. These data will allow a study of the optical properties = of an unbiased sample of AGN host galaxies. As detailes in 2.1, the = acquisition for the second iteration of this proposal failed. Otherwise, = the observations completed as scheduled, and no further problems were = reported. 1.5 Completed WF/PC-2 7016 (WF/PC-2 SM97 UV Monitor And Cool Down = Procedure) This operation was used as a post-SM97 cooldown of the WF/PC-2 CCD = to -88C with frequent monitoring of contamination via standard star = GRW+70D582 observations in filter F170W. An internal flat and a dark = calibration image were also taken. The observations were completed as = planed, and no problems were reported 1.6 Completed Two Sets of NIC 1/2/3 7054 (NICMOS Thermal Check on = COSTAR FOC Arm Deploy) The Near-Infrared Cameras 1/2/3 were used to make a thermal = check-out of the COSTAR FOC deployed optical bench arm. This is to = verify that the COSTAR FOC arm does not contribute to the thermal = background seen by NICMOS. The observations were executed as scheduled, = and no anomalies were noted. 1.7 Completed Two Sets of COSTAR/U FOC 7008 (COSTAR FOC M1/M2 Mirror = Arm Deployment) The COSTAR's FOC M1/M2 arms were returned to their pre-servicing = mission positions. The WF/PC-2's shutter was closed during and for 30 min= utes after the deployment of the FOC COSTAR arms. The FOC arm deployment = activities were executed with a combinations of real-time and stored = commanding. FOC M1/M2 arm deployments were not executed until the FOC = baseline observations were performed. No anomalies were reported. 1.8 Completed Two Sets of WF/PC-2 6837 (The Ultraviolet Properties of = "Dwarf" Seyfert 1 Nuclei) The WF/PC-2 was used to make ultraviolet observations of the = "dwarf", type one, Seyfert galaxies NGC-3884 and NGC-3227. This class of = "dwarf" Seyfert populate the faint end of the local AGN luminosity = function and hold important clues to the evolution of the entire AGN = population. The observations were completed as scheduled, and no = problems were reported. 1.9 Completed WF/PC-2 6363 (The Environments and Host Galaxies of = BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac object B2-1147+24 in order to study its environment and its host = galaxies. The goal is to investigate systematic trends in host galaxy = type, unusual morphological features, nearby companion galaxies, etc. = The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed FGS 6566 (Calibrating the Mass-Luminosity Relation at = the End of the Main Sequence) The FGSs were used to make astrometric observations of the = components of the low-mass, binary system, star GL-784 in order to make a = calibrations of the mass-luminosity relation at the end of the main = sequence. Observations were also made of several other reference stars. = The individual stars have masses of less than 0.2 solar masses. The = observations were completed as scheduled, and no problems were noted. 1.11 Completed Three Sets of WF/PC-2 7022 (WF/PC-2 SM97 Monitor) This calibration proposal is the routine internal monitor for = WF/PC-2, to be run twice weekly to monitor the health of the cameras. A = variety of internal exposures were obtained in order to provide a monitor = of the integrity of the CCD camera electronics in both bays (gain 7 and = gain 15), a test for quantum efficiency in the CCDs, an internal check on = the alignment of the WF/PC-2 optical chain, and a monitor for possible = buildup of contaminants on the CCD windows. This observation was = completed as scheduled, and no anomalies were noted. 1.12 Completed FGS 6495 (The Mass of the Bright Giant HD173764) The FGSs were used to make astrometric observations of bright, = spectrographic binary system, giant star HD173764 which has a 832 day = period. The objective is to determine the individual masses of the G4 = and B9 class pair. The observations were completed as scheduled, and no = anomalies were noted. 1.13 Completed FOC/96 7013 (FOC/96/FGS Alignment - Version 2) The Faint Object Camera (f/96) was used to make observations of = the star cluster M-35. This is to ensure that the V2-V3 position of the = FOC F/96 aperture has not been significantly altered by the Second = Servicing Mission activities. The observations were executed as planned, = and no problems were reported. 1.14 Completed WF/PC-2 6491 (The Optical Structure and Environments of = Z~1 3CR Quasars) The WF/PC-2 was used to make observations of radio galaxy/ quasar = 3C208 in order to study its optical structure and environments. This = should obtain better statistics on correlations seen in the Cycle 4 data, = such as the association of linear galaxies with other 3CR galaxies and = quasars. The observations were executed as scheduled, and no anomalies = were noted. 1.15 Completed WF/PC-2 6565 (Imaging of Proto-Planetary Nebulae) The WF/PC-2 was used to make high resolution observations of the = proto-planeatary nebula IRAS16594 which is known to show bipolar = morphologies in ground-based observations. The observations were executed= as scheduled, and no anomalies were noted. 1.16 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTULIR18. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.17 Completed WF/PC-2 6701 (Photometry of the Globular Cluster System = of the Sagittarius Dwarf Galaxy) The WF/PC-2 was used to make observations of the globular cluster = system Ter8 which is located the Sagittarius Dwarf Galaxy. Isochrone = fits to these data will yield the accurate age of this system. As = detailed in 2.1, both the acquisition and reacquisition for this proposal = failed. Otherwise, the observations completed as schedule, and no = further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 15 Successful : 13 As recorded in HSTAR 5990, the acquisition at 081/235044Z failed = to gyro control when the FGS-3 scan steps limit was exceeded. HSTAR 5991 details the acquistion and subsequent reacquisition = failures at 082/201600Z and 082/222203Z respectively. For each, the = search radius limit was exceeded. Scheduled Reacquisitions : 20 Successful : 19 2.2 FHST Updates: Scheduled : 26 Successful : 26 2.3 Operations Notes: Per HSTAR 5989, NICMOS, transitioned from boot to operate (in the = SAA) at 081/164000Z. However, the TPG1 clock stayed on and the detector = 1 temperature remained enabled. The NICMOS filter wheels were re-positioned at 081/1641Z. At 083/0144Z, the SSR was configured for continuous engineering = record. ESTR #2 remains powered on, but is not being used. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH